Minden Company manufactures a high-quality wooden birdhouse that sells for $20
per unit. Variable costs are $6 per unit, and fixed costs total $180,000. The
company sold 24,000 birdhouses to customers during 2018. The president of
Minden Company believes the following changes should be made in 2019:

1. the selling price of the birdhouse should be reduced by 10%

2. increase advertising by $30,000

Assume these changes are made to complete the following;

1) Calculate the number of units Minden Company must sell in 2019 in order to break-even.

2) Minden Company sells 26,000 units in 2019. Calculate the margin of safety reported by Minden Company in 2019.

3) Calculate the number of units Minden Company must sell in 2019 in order to earn a target profit equal to 25% of sales.

4) Calculate the number of units Minden Company must sell in 2019 in order to earn a net income that is 40% larger than the net income earned in 2018.
Answer #1

New sales price = $20 - (10%×$20) = $18

New fixed cost = $180000 + $30000 = $210000

New contribution margin per unit = $18 - $6 = $12

1)

Break even unit = fixed cost/contribution margin

= $210000/$12 = 17500 units

2)

Margin of safety = sales - break even sales

= 26000 - 17500

= 8500 units

3)

Let the number of units be X

Then contribution margin will be = 12X

And required profit = sales × 25% = 18X × 25% = 4.5X

Profit = contribution margin - fixed cost

4.54 = 12X - $210000

$210000 = 7.5X

X = units sold = $210000/7.5 = 28000 units

4)

Net income earned in 2018 =

Sales $480000
Less: Variable cost $144000
Contribution margin $336000
Fixed cost ($180000)
Net income $156000

Required net income = $156000 + ($156000×40%) = $218400

Contribution margin = fixed cost + net income

= $210000 + $218400

= $428400

Units to be sold = contribution margin / contribution margin per unit

= $428400/12

= 35700 units

I have given complete and detailed solution to your problem, in case of any query you can comment and

PLEASE DO GIVE POSITIVE RATING